A circular flower garden with a diameter 30 meters has a walkway around it. If the width of this walkway is 3 meters, what is the area?
Answer by checkley77(12844)

Area=pir^2
Area=3.14(30/2)^2
Area=3.14*15^2
Area=3.14*225
Area=706.5 m^2 is the area of the flower garden.
Area=3.14(33/2)^2
Area=3.14*16.5
Area=3.14*272.25
Area=854.865 m^2 is the area of the flower garden plus the walkway.
The diference in area is the area of the walkway.
854.865-706.5=148.365 m^2 is the area of the walkway.